OpenStudy (anonymous):

Please help! cos67degrees 30'= ?

OpenStudy (anonymous):

0.890

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

\[a. \sqrt{2-\sqrt{2}} \] \[b. \sqrt{2+\sqrt{2}}/2\] \[c. \sqrt{2+\sqrt{2}}\] \[d. \sqrt{2-\sqrt{2}}/2\] @muhammad9t5 Your answer doesn't match with the options :/

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

can you post screenshot of question?

OpenStudy (anonymous):

b. \[(\sqrt{2+\sqrt{2}})/2\] d. \[(\sqrt{2-\sqrt{2}})/2\]

OpenStudy (anonymous):

Sorry, I messed those two up

OpenStudy (anonymous):

\[67^{\circ} 30' = 67 + \frac{1}{2} \implies (\frac{135}{2})^{\circ}\]

OpenStudy (anonymous):

Is there only one cos in your question @IloveCharlie

OpenStudy (anonymous):

you can check that b is a nearly answer.

OpenStudy (anonymous):

Yes, @waterineyes, only one cos

OpenStudy (anonymous):

Can I get more help for this one? sin67degrees 30'= a., b., c., or d.?

OpenStudy (anonymous):

@muhammad9t5 @waterineyes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer 1.42?

OpenStudy (anonymous):

Can you please show me how you got that?

OpenStudy (anonymous):

yes sin67+30/60 0.920+1/2 0.920+0.5=1.42

OpenStudy (anonymous):

and that matches with? :/ sorry

OpenStudy (anonymous):

that's confusing me too.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Do you know what 1.42 matches with? a., b., c., or d.? I can't figure it out @waterineyes

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok... well... I guess I'll close the question then

OpenStudy (anonymous):

Thanks guys